Shift report 20010818 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Aug 18 2001 - 00:25:26 EDT

  • Next message: Ian Bearden: "Run 4264, understanding magnetic fields in BRAHMS"

    ------------------------------------------------------------
    Date: Sat Aug 18 00:25:26 2001, Shift: 16:00-24:00
    Supervisor: ykl, with: jjg
    Log book#7 pages 117 - 119
    
    Shift Summary:
    ------------------------------------------------------------
    1. Vacuume leak and other problems most of the time.
    Beam was cogged during the last one hour.
    2. Run 4264 has been in progress for an hour without
    D2 magnet on.
    3. T5 low voltage supply tripped once at the beginning of
    the shift. Anode3 of T3 cannot be turned on and has
    a leakage current of 75 microamp.
    
    D2 magnet as seen on the PET screen fluctuates +-20 amp.
    Experts from the control room says the current is steady
    when it is measured directory from the power supply.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Aug 18 2001 - 00:26:16 EDT